Would like the solution for current. resistance is divided by 100 so 5, 7.5, and 10011 R1 (500) A А A R2 (750 (2) E (10.0 V) 12 13 R3 (1000 (2) €2 (5.00 V)
Answer #1

putting the values of resistens 12 ST 5 1 2 = 5 512 - 105 I2 = 10 12 - 0.369 Amp Iz = 0.0777 Amp. l, = 12 + 5z = 0.369 + 0.09R = 5, R2 = 7.5, Rs = 100m I. Is VI2 I = I2 + Iz. R2 K, 10.0 in left side loop 5.o 10- Rili- R2F₂-s=o. RII + R2 [2 R (L2 + f₂